# MatCLIP: Light- and Shape-Insensitive Assignment of PBR Material Models

### v1

@@ Abstract.
Assigning realistic materials to 3D models remains a significant challenge in computer graphics. We propose MatCLIP, a novel method that extracts shape- and lighting-insensitive descriptors of Physically Based Rendering (PBR) materials to assign plausible textures to 3D objects based on images, such as the output of Latent Diffusion Models (LDMs) or photographs. Matching PBR materials to static images is challenging because the PBR representation captures the dynamic appearance of materials under varying viewing angles, shapes, and lighting conditions. By extending an Alpha-CLIP-based model on material renderings across diverse shapes and lighting, and encoding multiple viewing conditions for PBR materials, our approach generates descriptors that bridge the domains of PBR representations with photographs or renderings, including LDM outputs. This enables consistent material assignments without requiring explicit knowledge of material relationships between different parts of an object. MatCLIP achieves a top-1 classification accuracy of 76.6%, outperforming state-of-the-art methods such as PhotoShape and MatAtlas by over 15 percentage points on publicly available datasets. Our method can be used to construct material assignments for 3D shape datasets such as ShapeNet, 3DCoMPaT++, and Objaverse. All code and data will be released.
@@ 2. Related Work
Creating realistic materials for 3D models is a longstanding challenge in computer graphics, addressed through diverse approaches such as exemplar-based material assignment, RGB-only texturing, SVBRDF estimation, and diffusion-based material generation. Exemplar-based methods like PhotoShape (Park et al., 2018) classify material categories using photographs but lack robust embeddings for PBR materials, limiting generalization. Extensions like PhotoScene (Yeh et al., 2022) optimize material parameters through differentiable graphs but depend on proprietary Adobe Substances, reducing reproducibility. In contrast, our method uses contrastive learning to align embeddings across diverse lighting and geometric conditions, improving robustness and accessibility.
@@ 3.2. Material Representation
Method Top-1 Accuracy [%] $\uparrow$ Top-5 Accuracy [%] $\uparrow$
PhotoShape (ResNet-34) 57.21 77.39
PhotoShape (ResNet-50) 58.76 77.62
PhotoShape (ViT-B/16) 60.86 78.23
OpenCLIP (Max. Inscribed Square Crop, Max. Score to 42 Material Features) 16.13 31.49
OpenCLIP (Max. Inscribed Square Crop, Avg. Material Features) 19.90 35.34
MatAtlas (0.8 × LAB + 0.2 × CLIP, Max. Score to 42 Material Features) 18.77 36.79
MatAtlas (0.8 × LAB + 0.2 × CLIP, Avg. Material Features) 13.14 28.40
MatCLIP (Ours) 76.69 89.02
Table 1. Quantitative evaluation of our method. We compare to variants of PhotoShape (Park et al., 2018), simple crop matching using CLIP (Radford et al., 2021), and variants of the core method of MatAtlas (Ceylan et al., 2024). We either average or identify the max. score across the 42 material features of each of our materials in case of CLIP and MatAtlas. Accuracy is evaluated on Top-1 and Top-5 metrics, with higher values indicating better performance.
@@ 3.4. Sampling Textures from Stable Diffusion
To generate $I_{\text{target}}$, we condition Stable Diffusion on depth and edge information derived from the 3D model, ensuring that the generated textures align with the model’s geometry. Depth information is captured by rendering a single-view depth image, $I_{\text{depth}}$, from an informative perspective, as discussed in Section 3.2. In addition to depth conditioning, edge information is extracted using Canny edge detection on a diffuse gray rendering of the 3D model. This ensures fine geometric details, such as thin wires or fabric strips, are preserved in $I_{\text{target}}$. We also incorporate a CLIP text embedding derived from descriptive prompts suggested by ChatGPT (e.g., “a sturdy and durable chair with a robust design”), which guides the stylistic interpretation of the material. By combining depth, edge, and text conditioning, Stable Diffusion produces target images that capture both geometric accuracy and material diversity. Examples of depth images, edge maps, and resulting outputs are shown in Fig. 3.
@@ 4. Results
Our model is trained on a GPU cluster node with 4 NVIDIA A100 40GB cards for approximately 48 hours. Hyperparameters follow the contrastive loss strategy introduced in Alpha-CLIP (Sun et al., 2023). MatCLIP adapts this approach to align part-level embeddings with diverse material representations, as detailed in Section 3.2. This enables MatCLIP to robustly handle variations in shape and lighting conditions while maintaining consistent performance.
@@ 4.1. Classification Performance
We evaluate MatCLIP’s material classification performance by ranking materials in the test split of our dataset, using descriptors generated by the part encoder ($E_{\text{part}}$) and material encoder ($E_{\text{material}}$). Materials are ranked by cosine similarity between their embeddings. For each query shape from the test set, materials were matched from the entire set of available materials, without restricting classification to the substance categories defined for the 3DCoMPaT++ shapes. This ensured that the model was evaluated across the full database of materials. The Top-1 and Top-5 accuracies of our method, compared to PhotoShape (Park et al., 2018) and MatAtlas (Ceylan et al., 2024), are shown in Table 1.
For a fair comparison, we modified PhotoShape by replacing its original ResNet34 and ResNet50 backbones with the more recent ViT-B/16 architecture to better align with modern methods. This resulted in slightly improved performance of PhotoShape on our data. Additionally, we evaluated MatAtlas without restricting classification to the material classes suggested by ChatGPT, instead testing its ability to match materials from the full database. Although this adjustment reduced MatAtlas’s performance, our intent was to assess only the classification component, ensuring the comparison focused on embedding quality rather than substance category limitations. These adjustments aimed to standardize the evaluation setup and ensure all methods were tested under similar conditions. Despite these efforts to strengthen the baselines, MatCLIP achieves the highest accuracy, with a Top-1 accuracy of 76.69% and a Top-5 accuracy of 89.02%, significantly outperforming the next-best method, PhotoShape (ViT-B/16) which had a top-1 accuracy of 60.86%.
@@ 4.2. Generalization to Unseen Materials
Method Val.Acc. [%] $\uparrow$
PhotoShape (ResNet-34) 46.67
PhotoShape (ResNet-50) 49.07
PhotoShape (ViT-B/16) 44.27
OpenCLIP (Max. Score) 49.20
OpenCLIP (Avg. Features) 44.65
MatAtlas (0.8 × LAB + 0.2 × CLIP, Max.) 52.94
MatAtlas (0.8 × LAB + 0.2 × CLIP, Avg.) 48.93
MatCLIP (Ours) 55.47
Table 2. Model comparison using manually annotated ground-truth data. We evaluate the performance of assigning the correct substance derived from matched materials. We use the same models as in Table 1.
We then compared MatCLIP’s Top-1 ranked material assignments to those of competing methods, including PhotoShape (Park et al., 2018) and MatAtlas (Ceylan et al., 2024). The results are summarized in Table 2. Although MatCLIP consistently outperforms other approaches, the performance gap is narrower than in prior evaluations. This is likely due to the inherent ambiguity and subjectivity in manual annotations.
@@ 4.3. Ablation Study
To assess the impact of dataset diversity on material classification accuracy, we conducted an ablation study by varying the number of shapes and environment maps used to render materials. The full MatCLIP model relies on 42 renderings per material, generated using 6 distinct shapes and 7 HDR environment maps. We systematically reduced the dataset by testing subsets, including cases with a single shape across all environment maps (Plane 7 Env) and all shapes under a single environment map (All Shapes 1 Env). For each subset, we retrained MatCLIP and monitored Top-1 classification accuracy over the first 9,000 training steps.
The results, shown in Table 3, highlight the significant advantage of using the full dataset. The full model achieves the highest accuracy, significantly outperforming models trained on subsets. Notably, the worst performance was observed with a single shape under a single environment map (Plane 1 Env), emphasizing the importance of both geometric and lighting diversity for learning robust material representations.
@@
Plane Shape, 7 Environment Maps 53.19
@@
Table 3. Ablation study examining the effect of shape and environment map diversity on classification accuracy. The full model, which uses all six shapes and seven environment maps (42 renderings per material), performs best. Variants include: ’Plane 1 Env,’ using a single shape (plane) and one environment map; ’Plane 7 Env,’ using the plane with all seven environment maps; and ’All Shapes 1 Env,’ using all six shapes with a single environment map.
@@ 4.4. Qualitative Evaluations with Latent Diffusion Models
We conduct qualitative evaluations using outputs from Stable Diffusion (cf. Fig. 7) and Flux (cf. Fig. 8) to assess MatCLIP’s ability to generalize beyond the closed set of training materials. These evaluations focus on assigning materials to synthetic target images ($I_{\text{target}}$) generated using depth, edge, and text conditioning, as described in Section 3.2. Our model selects the best-matching material from the MatSynth database for each target image, producing visually coherent and contextually appropriate assignments, even for complex shapes and lighting conditions.
Figures 7 and 8 showcase examples where MatCLIP robustly matches materials to parts of a 3D shape. The visual results illustrate the effectiveness of our embeddings in achieving plausible and realistic material assignments across diverse target distributions.
@@ 4.5. Prompt-Based Design Exploration
In order to explore a larger design space, we leverage the power of ChatGPT to generate a large number of prompts. In general, we distinguish between two types of prompts. We use a regular prompt that usually only consists of a few words like ”a comfortable chair.” The prompts we try to obtain from ChatGPT we refer to as advanced prompts, since they usually consist of very detailed descriptions of materials and the relationship between the materials that are used for different parts of a shape. To allow ChatGPT to describe such relationships, we do not only use a textual prompt but also provide an exemplar image of the particular shape of interest (cf. Fig. 4). Using advanced prompts, we are able to obtain a larger variety than we were able to get using only regular prompts (cf. Fig. 9).
Refer to caption This image was generated by an LDM using the prompt ”A photograph of a very comfortable chair, high quality, white background.” Please suggest 3 different prompts to obtain more meaningful but very differently looking results that take the geometry into account. Here are three alternative prompts you can use to generate varied but geometrically related chairs using a latent diffusion model: (1) A futuristic armchair inspired by mid-century modern design, featuring metallic legs and a sleek seat. (2) A luxurious lounge chair with an organic design, curved lines, and upholstered with soft velvet fabric. (3) A modern Scandinavian-style chair with a minimalist wooden frame, light fabric cushions, and angled legs.
Figure 4. We exploit ChatGPT to obtain various prompts we then use to condition Flux. Top to bottom: The query target image, our query prompt, and the response from ChatGPT representing a set of advanced prompts we use to obtain the material variations shown in Fig. 9, middle row.
@@ 4.6. User Study
To evaluate the effectiveness of our method for assigning plausible materials to large non-textured shape collections, we conducted a user study on Amazon Mechanical Turk (MTurk). Specifically, we selected 100 chair shapes at random from the 3DCoMPaT++ dataset and used Stable Diffusion, as detailed in Section 3.4, to generate plausible material assignments. For each chair shape, we created approximately five different sets of material assignments, resulting in a total of 506 textured chair shapes.
To benchmark our approach, we compared it to the procedural material assignment method from 3DCoMPaT++, which procedurally assigns materials to shape parts based on predefined substance categories. All chair shapes, including those textured using our method and the procedural approach, were rendered using Blender’s physically-based path tracer. To ensure a fair comparison, we used identical rendering parameters for all models, including the same camera settings (intrinsic and extrinsic), environment map, and overall scene composition.
We constructed a dataset of image pairs, each consisting of one chair textured using our method and one textured using the procedural approach. The order of images within each pair was randomized, and participants on MTurk were asked to choose the chair they believed had the more plausible material assignment (cf. Fig. 6). We collected two independent responses (HITs) per image pair to increase reliability, resulting in 1,012 ranked decisions.
@@ 5. Limitations
Generalization to unseen material distributions remains an open challenge. While qualitative evaluations (cf. Section 4.4) suggest that MatCLIP can produce visually coherent material assignments for unseen scenarios, quantitative evaluations using manually annotated substances show room for improvement. The Top-1 accuracy drops from 76.69% on the validation set of known materials (cf. Table 1) to 55.47% for manual annotations (cf. Table 2). However, this reduction is not unique to our method; other methods also experience performance drops under similar conditions. We believe this is mainly due to the uncertainty in the ground truth annotations, as the task of material assignment from RGB images is also very difficult for human labelers.
Figure 5. Our model struggles in cases where Stable Diffusion adds additional geometry (e.g., wrinkles on the left side) or where no material from the material database would give a good match (right side).

### v2

@@ Abstract.
Assigning realistic materials to 3D models remains a significant challenge in computer graphics. We propose MatCLIP, a novel method that extracts shape- and lighting-insensitive descriptors of Physically Based Rendering (PBR) materials to assign plausible textures to 3D objects based on images, such as the output of Latent Diffusion Models (LDMs) or photographs. Matching PBR materials to static images is challenging because the PBR representation captures the dynamic appearance of materials under varying viewing angles, shapes, and lighting conditions. By extending an Alpha-CLIP-based model on material renderings across diverse shapes and lighting, and encoding multiple viewing conditions for PBR materials, our approach generates descriptors that bridge the domains of PBR representations with photographs or renderings, including LDM outputs. This enables consistent material assignments without requiring explicit knowledge of material relationships between different parts of an object. MatCLIP achieves a top-1 classification accuracy of 76.6%, outperforming state-of-the-art methods such as PhotoShape and MatAtlas by over 15 percentage points on publicly available datasets. Our method can be used to construct material assignments for 3D shape datasets such as ShapeNet, 3DCoMPaT++, and Objaverse. All code and data will be released at https://birsakm.github.io/matclip/.
@@ 2. Related Work
Creating realistic materials for 3D models is a longstanding challenge in computer graphics, addressed through diverse approaches such as exemplar-based material assignment, RGB-only texturing, SVBRDF estimation, and diffusion-based material generation. Methods based on exemplars like PhotoShape (Park et al., 2018) classify material categories using photographs but lack robust embeddings for PBR materials, limiting generalization. Extensions like PhotoScene (Yeh et al., 2022) optimize material parameters through differentiable graphs. The best-matching material graph is identified by nearest neighbor search using a VGG feature distance. Similarly, PSDR-Room (Yan et al., 2023) and MaPa (Zhang et al., 2024) utilize CLIP in a zero-shot manner between pre-rendered thumbnails of material graphs and image crops to obtain best-matching material graphs. These methods depend on the proprietary Adobe Substances material library, limiting external reproducibility by the community. The Visual Language of Fabrics (VLF) (Deschaintre et al., 2023) fine-tunes CLIP on a curated dataset of (rendering, description) pairs exclusively for fabrics to improve the quality of text-based and image-based retrieval. In contrast, our method uses contrastive learning to align embeddings across diverse lighting and geometric conditions, improving robustness and accessibility. We want to emphasize that the idea of doing CLIP-based material retrieval has also been adopted by the methods cited above. The novelty of our approach is that we find an embedding for materials that is robust to geometry and lighting conditions but still within a CLIP space.
@@ 3.2. Material Representation
Table 1. Combined quantitative evaluation of all methods across four conditions: (1) main evaluation, and generalization to unseen (2) shape categories, (3) lighting, and (4) materials. We compare to variants of PhotoShape (Park et al., 2018), simple crop matching using CLIP (Radford et al., 2021), and variants of the core method of MatAtlas (Ceylan et al., 2024). For CLIP and MatAtlas, accuracy is evaluated based on Top-1 (T-1) and Top-5 (T-5) metrics by taking the max. score (v1) or averaging (v2) across 42 material features. Higher accuracy values indicate better performance. Note that the values in ‘Unseen Materials’ refer to substance-only assignment on a manually annotated dataset of LDM outputs and are therefore not directly comparable.
Method Main Evaluation Unseen Shapes Unseen Lighting Unseen Materials
T-1 [%] $\uparrow$ T-5 [%] $\uparrow$ T-1 [%] $\uparrow$ T-5 [%] $\uparrow$ T-1 [%] $\uparrow$ T-5 [%] $\uparrow$ T-1 [%] $\uparrow$
PhotoShape (ResNet-34) 57.21 77.39 36.63 59.19 18.83 35.19 46.67
PhotoShape (ResNet-50) 58.76 77.62 38.71 61.42 20.67 37.31 49.07
PhotoShape (ViT-B/16) 60.86 78.23 40.42 62.22 19.70 35.41 44.27
OpenCLIP (v1) 16.13 31.49 13.56 26.15 12.63 22.89 49.20
OpenCLIP (v2) 19.90 35.34 16.07 29.28 13.70 24.28 44.65
MatAtlas (v1) 18.77 36.79 14.20 28.50 11.97 23.43 52.94
MatAtlas (v2) 13.14 28.40 9.03 20.93 6.62 14.71 48.93
MatCLIP (Ours) 76.69 89.02 54.74 73.22 35.31 51.90 55.47
@@ 3.4. Sampling Textures from Stable Diffusion
To generate $I_{\text{target}}$, we condition Stable Diffusion on depth and edge information derived from the 3D model, ensuring that the generated textures align with the model’s geometry. Depth information is captured by rendering a single-view depth image, $I_{\text{depth}}$, from an informative perspective, as discussed in Section 3.2. In addition to depth conditioning, edge information is extracted using Canny edge detection on a diffuse gray rendering of the 3D model. This ensures fine geometric details, such as thin wires or fabric strips, are preserved in $I_{\text{target}}$. We also incorporate a CLIP text embedding derived from descriptive prompts suggested by ChatGPT (e.g., “a sturdy and durable chair with a robust design”), which guides the stylistic interpretation of the material. By combining depth, edges, and text conditioning, Stable Diffusion produces target images that capture both geometric accuracy and material diversity. Examples of depth images, edges, and resulting outputs are shown in Fig. 3.
@@ 4. Results
We base our model on a pre-trained ViT-B/16 version of Alpha-CLIP. In particular, our part encoder $E_{part}$ is based on its image encoder, taking an RGBA image as input. For the material encoder $E_{material}$, we leverage Alpha-CLIP’s fine-tuned transformer-based text encoder and use a linear layer to map to the transformer’s input dimension. For the training procedure, we use a batch size of 256 and a learning rate of 1.0e-4. Our model is trained on a GPU cluster node with 4 NVIDIA A100 40GB cards for approximately 48 hours. Hyperparameters follow the contrastive loss strategy introduced in Alpha-CLIP (Sun et al., 2023). MatCLIP adapts this approach to align part-level embeddings with diverse material representations, as detailed in Section 3.2. This enables MatCLIP to robustly handle variations in shape and lighting conditions while maintaining consistent performance.
@@ 4.1. Classification Performance
We evaluate MatCLIP’s material classification performance by ranking materials in the test split of our dataset, using descriptors generated by the part encoder ($E_{\text{part}}$) and material encoder ($E_{\text{material}}$). Materials are ranked by cosine similarity between their embeddings. For each query shape from the test set, materials were matched from the entire set of available materials, without restricting classification to the substance categories defined for the 3DCoMPaT++ shapes. This ensured that the model was evaluated across the full database of materials. The Top-1 and Top-5 accuracies of our method, compared to PhotoShape (Park et al., 2018) and MatAtlas (Ceylan et al., 2024), are shown in Table 1 (Main Evaluation).
For a fair comparison, we modified PhotoShape by replacing its original ResNet34 and ResNet50 backbones with the more recent ViT-B/16 architecture to better align with modern methods. This resulted in slightly improved performance of PhotoShape on our data. Additionally, we evaluated MatAtlas without restricting classification to the material classes suggested by ChatGPT, instead of testing its ability to match materials from the full database. Although this adjustment reduced MatAtlas’s performance, we intended to assess only the classification component, ensuring the comparison focused on embedding quality rather than substance category limitations. These adjustments aimed to standardize the evaluation setup and ensure all methods were tested under similar conditions. Despite these efforts to strengthen the baselines, MatCLIP achieves the highest accuracy, with a Top-1 accuracy of 76.69% and a Top-5 accuracy of 89.02%, significantly outperforming the next-best method, PhotoShape (ViT-B/16) which had a top-1 accuracy of 60.86%.
@@
4.2. Generalization to Unseen Shape Categories
While our network was trained only on renderings of chairs from the 3DCoMPaT++ shape dataset (Li et al., 2022; Slim et al., 2023), it is not restricted to chairs but generalizes well to arbitrary shape categories. For our experiments, we rendered 3007 images across all non-chair shape categories of 3DCoMPaT++. We removed very bad renderings showing heavy over- or underexposure to finally obtain 2997 renderings and corresponding 11k part-mask pairs to run the actual experiments. As shown in Table 1 (Unseen Shapes), our results on unseen shape categories show a slight performance drop, our method however still performs significantly better than the baseline methods. While we used only chairs for training for this proof of concept, we believe that the performance of our method could be improved by incorporating a larger variety of different shape categories into the training procedure.
4.3. Generalization to Unseen Lighting Conditions
We focus on studio lighting conditions to train our network because we argue that this best reflects the lighting conditions of product and shape images generated by latent diffusion models (LDMs) like Stable Diffusion (Rombach et al., 2021) and Flux (Labs, 2023). Still, our method is able to generalize well to unseen and also more extreme (e.g. colored) lighting as we show in Table 1 (Unseen Lighting). For our experiment, we download 769 HDRI environment maps from all available categories (e.g. outside, inside, daylight, artificial lighting, colored lighting) from Polyhaven and render 4502 images. Again, we remove outliers showing heavy over- or underexposure to obtain 4475 images and 18k part-mask pairs. While the performance of our model both on the Top-1 and Top-5 accuracy is lower than on unseen shapes, it still performs significantly better than the baseline methods we compare to. We believe that there is still room for improvement by incorporating a larger number of different lighting scenarios (e.g. natural lighting captured at different day times, colored lighting, single- and multi-source lights) into our training dataset.
@@ 4.4. Generalization to Unseen Materials
We then compared MatCLIP’s Top-1 ranked material assignments to those of competing methods, including PhotoShape (Park et al., 2018) and MatAtlas (Ceylan et al., 2024). The results are summarized in Table 1 (Unseen Materials). Although MatCLIP consistently outperforms other approaches, the performance gap is narrower than in prior evaluations. This is likely due to the inherent ambiguity and subjectivity in manual annotations.
@@ 4.5. Ablation Study
To assess the impact of dataset diversity on material classification accuracy, we conducted an ablation study by varying the number of shapes and environment maps used to render materials. Our full model relies on 42 renderings per material, generated using 6 distinct shapes and 7 HDR environment maps. We systematically reduced the dataset by testing subsets, including cases with a single shape across all environment maps and all shapes under a single environment map. For each subset, we retrained MatCLIP and monitored Top-1 classification accuracy over the first 9,000 training steps.
The results highlight the significant advantage of using the full dataset (cf. Table 2). The full model achieves the highest accuracy, significantly outperforming models trained on subsets. Notably, the worst performance was observed with a single shape under a single environment map, emphasizing the importance of both geometric and lighting diversity for learning robust material representations.
@@
Table 2. Ablation study examining the effect of shape and environment map diversity on classification accuracy. The full model, which uses all 6 shapes and 7 environment maps (42 renderings per material), performs best.
@@
Plane Shape, 3 Environment Maps 49.53
Plane Shape, 5 Environment Maps 52.47
Plane Shape, All Environment Maps 53.19
3 Shapes, 1 Environment Map 52.24
5 Shapes, 1 Environment Map 54.54
@@ 4.6. Qualitative Evaluations with Latent Diffusion Models
We conduct qualitative evaluations using outputs from Stable Diffusion (cf. Figure 5) and Flux (cf. Figures 6 and 7) to assess MatCLIP’s ability to generalize beyond the closed set of training materials. These evaluations focus on assigning materials to synthetic target images ($I_{\text{target}}$) generated using depth, edge, and text conditioning, as described in Section 3.2. Our model selects the best-matching material from the MatSynth database for each target image, producing visually coherent and contextually appropriate assignments, even for complex shapes and lighting conditions.
Figures 5, 6, and 7 showcase examples where MatCLIP robustly matches materials to parts of a 3D shape. The visual results illustrate the effectiveness of our embeddings in achieving plausible and realistic material assignments across diverse target distributions. Please refer to the supplementary materials for more results.
@@ 4.7. Prompt-Based Design Exploration
To explore a larger design space, we leverage the power of ChatGPT to generate a large number of prompts. In general, we distinguish between two types. We use a regular prompt that usually only consists of a few words like ”a comfortable chair.” The prompts we try to obtain from ChatGPT we refer to as advanced prompts, since they usually consist of very detailed descriptions of materials and the relationship between the materials that are used for different parts of a shape. To allow ChatGPT to describe such relationships, we do not only use a textual prompt but also provide an exemplar image of the shape of interest. Using advanced prompts, we are able to obtain a larger variety than we were able to get using only regular prompts (cf. Figures 1 and 4 in the supplementaries).
@@ 4.8. User Study
To evaluate the effectiveness of our method for assigning plausible materials to large non-textured shape collections, we conducted a user study on MTurk. Specifically, we selected 100 chair shapes at random from 3DCoMPaT++ and used Stable Diffusion, as detailed in Sec. 3.4, to generate plausible material assignments. For each chair shape, we created approximately five different sets of material assignments, resulting in a total of 506 textured chair shapes.
To benchmark our approach, we compared it to the procedural material assignment method from 3DCoMPaT++, which procedurally assigns materials to shape parts based on predefined substance categories. All chair shapes, including those textured using our method and the procedural approach, were rendered using Blender. To ensure a fair comparison, we used identical rendering parameters for all models, including the same camera settings (intrinsic and extrinsic), environment map, and overall scene composition.
We constructed a dataset of image pairs, each consisting of one chair textured using our method and one textured using the procedural approach. The order of images within each pair was randomized, and participants on MTurk were asked to choose the chair they believed had the more plausible material assignment. We collected two independent responses (HITs) per image pair to increase reliability, resulting in 1,012 ranked decisions.
@@ 5. Limitations and Future Work
Generalization to unseen material distributions remains an open challenge. While qualitative evaluations (cf. Section 4.6) suggest that MatCLIP can produce visually coherent material assignments for unseen scenarios, quantitative evaluations using manually annotated substances show room for improvement. The Top-1 accuracy drops from 76.69% on the validation set of known materials (cf. Table 1 (Main Evaluation)) to 55.47% for manual annotations (cf. Table 1, (Unseen Materials)). However, this reduction is not unique to our method; other methods also experience performance drops under similar conditions. We believe this is mainly due to the uncertainty in the ground truth annotations, as the task of material assignment from RGB images is also very difficult for human labelers.
In the future, we want to investigate potential alternative use cases like using our model in classifier-guided diffusion-based material optimization instead of just retrieval.
Figure 4. MatCLIP struggles when Stable Diffusion (left side) or Flux (right side) add additional geometry (e.g., wrinkles on the left side) or when no material from the material database would give a good match (right side).